All Systems Operational
Core Operational
90 days ago
100.0 % uptime
Today
S1 Operational
90 days ago
99.99 % uptime
Today
S2 Operational
90 days ago
100.0 % uptime
Today
S3 Operational
90 days ago
99.99 % uptime
Today
S6 Operational
90 days ago
99.7 % uptime
Today
S10 Operational
90 days ago
95.51 % uptime
Today
S11 Operational
90 days ago
96.36 % uptime
Today
S35 Operational
90 days ago
99.94 % uptime
Today
S33 Operational
90 days ago
99.99 % uptime
Today
S80 Operational
90 days ago
99.88 % uptime
Today
NG17 Operational
90 days ago
100.0 % uptime
Today
TN14 Operational
90 days ago
100.0 % uptime
Today
Phone Services (VOIP) Operational
90 days ago
100.0 % uptime
Today
Residential Phone Services Operational
90 days ago
100.0 % uptime
Today
Business Phone Services Operational
90 days ago
100.0 % uptime
Today
Carrier / Transit Providers ? Operational
90 days ago
100.0 % uptime
Today
Cogent Operational
90 days ago
100.0 % uptime
Today
CityFibre Operational
90 days ago
100.0 % uptime
Today
SSE Operational
90 days ago
100.0 % uptime
Today
BT Wholesale Operational
90 days ago
100.0 % uptime
Today
Lumen Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Past Incidents
Sep 22, 2021

No incidents reported today.

Sep 21, 2021
Resolved - All work has now been completed and customers should be coming back online. Your equipment may need a reboot, however if service has not been restored please get in touch with our support team on 0114 299 1699 or email support@pinemedia.net. Thank you for your patience.
Sep 21, 11:52 BST
Update - Our engineers are carrying out the final section of work required to restore service.

This should be completed in the next 30 - 50 Mins.

Thank you for your patience.
Sep 21, 11:19 BST
Update - The final replacement cable is now in place, our fibre splicers are working on commissioning it as the last activity required, we hope to have all circuits restored by 11:00 tomorrow.
Sep 20, 18:53 BST
Update - Our engineers are still working on getting the remaining customers back online. The work required has proven to be much more complex and difficult to complete but please be assured we are working on this as an utmost priority.
Sep 20, 17:13 BST
Update - Our engineers are currently re-cabling the road crossing which serves Broomgrove & Southgrove road
Sep 20, 09:09 BST
Update - Service has now been restored for the majority of circuits, further work is required for some customers on Broomgrove & Southgrove Road. This work is more complex as a road crossing is involved, we expect these circuits to be restored by 11:00 on Monday
Sep 18, 11:08 BST
Update - Splicing work has been completed but we are still seeing some circuits down, a secodary issue has been identified and our engineers are working to resolve it. We expect the majority of circuits to be restored in this area by 11:00 tomorrow, we sincerely apologise for the inconvenience casued - a case has been opened with BT regarding the damage they have caused.
Sep 17, 19:50 BST
Update - Our engineers have run new fibre cable to replace one of the damaged cables and are currently splicing the ends back together as we speak. We expect some customers to be coming back online slowly as this process is completed.
There may be further delays to some other customers due to the route that one of the other damaged fibre cables takes making replacing it much harder. We will update this page as soon as we know more.
Sep 17, 15:48 BST
Update - Our engineers are continuing to work on this issue.
Sep 17, 14:06 BST
Identified - We have identified a fibre break on Ecclesall Road caused by BT contractors working in the vicintity. We have engineers on site preparing to replace the damaged section of cable. We expect service to be restored before 16:00 today.
Sep 17, 11:16 BST
Update - We suspect a fibre break, our engineers are on-site at the location identified by our diagnostics
Sep 17, 11:15 BST
Update - We are continuing to investigate this issue.
Sep 17, 11:05 BST
Investigating - We are aware of a connectivity issue affecting FTTP customers and are currently investigating the issue.
Sep 17, 10:46 BST
Resolved - One of our datacentres lost power at around 03:00, we were able to restore with backup power by 07:00. No further downtime is expected.
Sep 21, 03:00 BST
Sep 20, 2021
Resolved - This incident has been resolved.
Sep 20, 16:03 BST
Monitoring - A fix has been implemented and we are monitoring the results.
Sep 20, 16:01 BST
Investigating - We are currently investigating this issue.
Sep 20, 15:57 BST
Sep 19, 2021

No incidents reported.

Sep 18, 2021
Sep 17, 2021
Sep 16, 2021
Resolved - This incident has been resolved.
Sep 16, 16:30 BST
Investigating - A service outage was identified that caused some customers to be offline. We are currently investigating this issue.
Sep 16, 16:20 BST
Sep 15, 2021

No incidents reported.

Sep 14, 2021

No incidents reported.

Sep 13, 2021

No incidents reported.

Sep 12, 2021

No incidents reported.

Sep 11, 2021

No incidents reported.

Sep 10, 2021

No incidents reported.

Sep 9, 2021

No incidents reported.

Sep 8, 2021

No incidents reported.